In a population of interest, 1 in 2000 women are affected with haemophilia A, an X-linked recessive condition.

a) What is the carrier frequency in this population for Haemophilia A?

b) What is the expected frequency of men affected with Haemophilia A?
Answer #1

A) Genotypic frequentcy of disease female is 1/2000=0.0005

The allelic frequency of disease allele (Xh) is √0.0005= 0.0223606798 or 0.022

The allelic frequency of XH is 1- Xh = 1-0.22= 0.978

Carrier frequency is 2×XH× Xh = 2×0.978×0.022= 0.043032

B) The expected frequency of affected male is Xh= 0.0223606798 or 0.022 on round off upto three decimal place.

Because the genotype of affected male has XhY. Affected male has one Xh allele only.